# Event-Based Control via Sparsity-Promoting Regularization: A Rollout Approach with Performance Guarantees

### v2

@@ I Introduction
Sparse control, in which the control signal remains zero for as long as possible within a control horizon, has attracted much attention in systems and control. Such control methods facilitate achieving control objectives with limited resource usage in networked control systems, since sparse signals admit compact representations [1]. Moreover, sparse control has the potential to reduce power consumption by saving actuation durations, which is crucial in applications such as railways [2] and electric vehicles [3]. These advantages have motivated development on the theoretical foundations of sparse control, including controllability [4], stability [5], and optimal control [6, 7]. A central challenge lies in developing frameworks that achieve desirable control performance with sparse inputs, due to the fundamental trade-off between closed-loop performance and input sparsity [8].
Sparsity of the control signal can be characterized in various ways, such as by the duration of nonzero inputs in an element-wise sense or by the number of active actuators at each time step. Among these notions of sparsity, this paper focuses on reducing the actuation rate, which gives rise to intermittent actuation, where the control input remains identically zero over consecutive time intervals. Such sparsity arises in practice; for example, in energy packet networks [9], where an on-demand power supply delivers energy to actuators intermittently. Motivated by this practical relevance, we investigate a sparse control method that achieves desirable control performance under intermittent actuation.
@@
In line with this direction, controller design methods for achieving desired performance under intermittent actuation have been investigated. For example, [10, 11, 12] have examined optimal control problems under constraints on the number of control actions. Alternatively, [13] and [14] have incorporated sparsity-promoting regularization terms into the optimization problem instead of enforcing explicit constraints. These studies provide near-optimal actuation strategies with theoretical performance bounds; however, their problem settings are restricted, such as considering noise-free systems [14, 12] or excluding control energy from the objective function [10, 11, 13]. For the case where the objective function consists of the state trajectory and the control input, [15] studied quadratic cost minimization for noise-affected systems under a threshold-based event-triggering rule combined with packetized dead-beat control. While their work quantifies the trade-off between communication and control performance, the actuation timings are determined by the predesigned threshold rule rather than being optimized. This challenge highlights the need for a principled method that jointly optimizes discrete actuation timings and continuous control laws to balance control performance and actuation rate, while providing provable performance improvements.
@@
One promising approach to such co-design problems is the rollout algorithm, a sequential optimization technique in dynamic programming [16, 17]. In a rollout scheme, control laws and scheduling decisions over a finite horizon are optimized sequentially. For example, [18] has addressed quadratic cost minimization under an average communication rate constraint and developed a rollout-based event-triggered control method that provides both performance and stability guarantees. This method achieves strictly better performance than the periodic control strategy under the same communication rate; however, its flexibility is limited by the assumption of a uniform communication rate within each horizon. Furthermore, [19] has studied the simultaneous minimization of a finite-horizon quadratic cost and the number of control actions. The authors proposed an algorithm that sequentially minimizes the one-step cost involving the control input and actuation timing, providing theoretical performance guarantees but no stability results.
In this paper, we propose a controller design method that aims to balance control performance and actuation rate, which provides both performance and stability guarantees. To promote sparsity in the control signal, we introduce an actuation-rate regularization term into the optimization problem, instead of imposing a constraint on the number of control actions. We formulate the problem as a co-design of control inputs and actuation timings, and seek to find control and triggering policies that minimize an objective function combining the LQ cost with a sparsity-promoting term. Based on the rollout strategy, control and triggering policies over multiple stages are optimized, assuming that periodic control, where the input is applied at fixed time instants, is used beyond the horizon. This procedure provides a more general framework than that in [19]. We show that the proposed method achieves theoretical performance bounds with respect to periodic control and ensures mean-square stability of the closed-loop system. While our techniques in performance analysis are similar to those in [18], it is not straightforward to apply their results to the present work, since the trade-off between control performance and communication rate is formulated in a different way. Moreover, whereas [18] establishes stochastic stability of the system in terms of ergodicity, obtained as a byproduct of the performance analysis, this stability notion cannot directly extend to the output-feedback setting considered in this paper. A numerical example demonstrates the effectiveness of the proposed algorithm.
@@
The remainder of this paper is organized as follows. In section II, we describe the problem formulation. Section III provides some preliminaries and presents the proposed algorithm, followed by a theoretical analysis of our framework. Section V illustrates a numerical example, and Section VI concludes this paper.
@@ III Proposed Rollout-Based algorithm
This section presents algorithm design using a rollout approach. In Section III-A, we give preliminaries on the rollout method. Section III-B derives the optimal periodic policy that is used in the algorithm design, and section III-C describes the algorithm design.
@@ III-C Algorithm Design
By the tower property of conditional expectation [22, Th. 5.1.6], we have
@@ IV Performance and Stability Guarantees of the Proposed Algorithm
Assumption 2(i) ensures that the steady-state Kalman gain $G$ has full row rank, since the covariance matrix $\tilde{\Sigma}$ is positive definite. Assumption 2(ii) guarantees that the minimum eigenvalue of $Q$ is strictly positive [23, Lemma 10.4.1]. When Assumption 2(iii) holds, $\Omega_{x_{0}}$ is positive definite since $(A,\Omega_{w}^{1/2})$ is controllable. Furthermore, this assumption implies $G_{k}=G$ for all $k$, i.e., the Kalman filter is stationary, which is the case of interest in this study.
@@
In (66), we write $\beta^{(m)}$ instead of $\beta^{(m)}_{k}$, since $\beta^{(m)}_{k}$ is independent of the time index $k$ under Assumption 2. By construction of Algorithm 1, we have $f_{m}(x)\geq 0$ for all $x\in\mathbb{R}^{n_{x}}$. Then, it follows that the Markov chain $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}_{\geq 0}}$ is ergodic, i.e., positive Harris recurrent [24, Ch. 14].
@@ VI Conclusion
We have proposed a co-design method of control and triggering policies that balances the trade-off between control performance and input sparsity. To this end, we formulated the optimal control problem with actuation-rate regularization and developed a rollout-based algorithm using a receding horizon approach. The proposed algorithm provides a theoretical performance guarantee and ensures mean-square stability of the closed-loop system. Its effectiveness has been demonstrated through a numerical simulation. Future work includes establishing performance guarantees for the discounted-cost problem and deriving strict performance improvement results over the periodic control strategy.
@@
Assumption 1 and Lemma 3 ensure that $\tilde{P}_{p}$ is positive semidefinite at $\alpha=1$ [21, Sec. 8.4]. Assume that there exists a nonzero vector $\xi\in\mathbb{R}^{n_{x}}$ such that $\xi^{\top}\tilde{P}_{p}\xi=0$. Then, we consider the following optimal control problem for a noise-free system with the initial condition $x_{0}=\xi$:
@@
To establish the coerciveness of $V_{1}$, it is sufficient to show that the sublevel sets $\mathcal{L}_{r}(x)\coloneqq\{x\in\mathbb{R}^{n_{x}}\colon V_{1}(x)\leq r\}$ are precompact, i.e., their closures are compact, for each $r\in\mathbb{R}_{>0}$ [24, Sec. 9.4]. Since $P^{(1)}_{0}$ is positive definite by Lemmas 4 and 5, we have
@@
The set on the right-hand side is bounded for every $r\in\mathbb{R}_{>0}$, and hence $\mathcal{L}_{r}(x)$ is also bounded. Moreover, the continuity of $V_{1}$ implies that $\mathcal{L}_{r}(x)=V_{1}^{-1}([0,r])$ is closed [27, Th. 4.8]. By the Heine–Borel theorem, $\mathcal{L}_{r}(x)$ is compact [27, Th. 2.41], and thus its closure is also compact for each $r\in\mathbb{R}_{>0}$ [27, Th. 2.27], which establishes the claim. ∎
@@
Conditioning on $\hat{\chi}_{0}$ and applying both the tower property and the linearity of conditional expectation [22, Th. 5.1.2], the left-hand side of (125) becomes
@@
where $\pi_{\hat{\chi}}$ is a unique invariant measure of the chain $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}_{\geq 0}}$ [24, Th. 14.0.1]. Note that the existence of $\pi_{\hat{\chi}}$ is guaranteed by Theorem 2. The above integrand is strictly positive due to the open set irreducibility of $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}_{\geq 0}}$ (see Appendix -F for details). Applying the Stolz–Cesàro theorem [29] to (143) yields
@@
In the following, we show that $P$ is a transition probability kernel and that the process $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}_{\geq 0}}$ has the Markov property [30, Def. 8.4.3], i.e.,
@@
where $g_{m}(x)\coloneqq x^{\top}P^{(m)}_{0}x+\operatorname{tr}(P^{(m)}_{0}\Sigma)+\beta^{(m)}+\gamma^{(m)}$ and $h_{i,m}(x)\coloneqq g_{i}(x)-g_{m}(x)$. This implies that $\mathcal{C}_{m}$ is Borel measurable since each inverse image is open and Borel measurable [27, Th. 4.8]. Therefore, for any $m\in\mathcal{M}$, we have $\varphi^{-1}(\{m\})=\mathcal{C}_{m}\in\mathscr{B}(\mathbb{R}^{n_{x}})$, which establishes the measurability of $\varphi$ [31, Th. 1.12].
@@
Then, $\hat{\chi}_{\ell+1}=\Psi(\hat{\chi}_{\ell},\bar{\omega}_{\ell})$ for a given $\varphi$. For each $x$ and $\omega$, the mapping $(x,\omega)\mapsto\Phi_{m}x+\Gamma_{m}\omega$ is continuous, and hence measurable [31, Th. 1.7]. Since the indicator function is measurable, the mapping $(x,\omega)\mapsto\mathbb{I}_{\mathcal{C}_{m}}(x)(\Phi_{m}x+\Gamma_{m}\omega)$ is measurable. Therefore, $\Psi$ is also measurable for any $m\in\mathcal{M}$ [31, Th. 1.8].
@@
for later use. By construction, $\bar{\omega}_{\ell}$ is a measurable function of $(e_{\ell h},\,\bar{\mathcal{W}}_{\ell})$, and hence $\sigma(\bar{\omega}_{\ell})\subseteq\sigma(e_{\ell h})\vee\sigma(\bar{\mathcal{W}}_{\ell})$ holds from [31, Th. 1.12]. Moreover, the independence of the noise processes yields $\sigma(\bar{\mathcal{W}}_{\ell})\perp\sigma(\mathcal{I}_{\ell h})$ [32, Lemma 4.6]. Since $e_{\ell h}=x_{\ell h}-\mathbb{E}[x_{\ell h}\mid\mathcal{I}_{\ell h}]$ under Gaussian noises implies that $e_{\ell h}$ is independent of $\mathcal{I}_{\ell h}$ [32, Lemma 14.1], we have
@@
which implies $\sigma(\bar{\omega}_{\ell})\perp\sigma(\mathcal{I}_{\ell h})$ [32, Cor. 4.7]. Since $\mathscr{F}_{\ell}\subseteq\sigma(\mathcal{I}_{\ell h})$, we obtain
@@
by [22, Ex. 5.1.5], where $\nu_{\bar{\omega}}$ denotes the distribution of $\bar{\omega}_{\ell}$. For a given $x\in\mathbb{R}^{n_{x}}$, combining (155), (164), and (166) yields
@@
from (166) and (168) [22, Th. 5.1.5].
For a fixed $x\in\mathbb{R}^{n_{x}}$, we define the mapping $\Psi_{x}\colon\omega\mapsto\Psi(x,\omega)$. We note that $\Psi_{x}$ is measurable for each $x$ since $\Psi$ is measurable on $\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{x}h}$ [31, Th. 8.5]. Then, for every $\mathcal{A}\in\mathscr{B}(\mathbb{R}^{n_{x}})$, it holds that
@@
which means that the mapping $\mathcal{A}\mapsto P(x,\mathcal{A})$ is a pushforward measure induced by $\Psi_{x}$; hence, $\mathcal{A}\mapsto P(x,\mathcal{A})$ is a probability measure on $\mathscr{B}(\mathbb{R}^{n_{x}})$ [32, Lemma 1.24]. Furthermore, $x\mapsto P(x,\mathcal{A})$ is measurable for a fixed $\mathcal{A}$, which follows by combining (167) and Fubini’s theorem. Therefore, $P(x,\mathcal{A})$ satisfies the definition of the transition probability kernel [24, Sec. 3.4].
By the property of conditional expectation [30, Th. 5.4.6], for any sub $\sigma$-algebra $\mathscr{G}\subseteq\mathscr{B}(\mathbb{R}^{n_{x}})$, we have
@@
Since $\Omega_{v}\succ 0$ and Assumption 2 holds, the matrix $\Omega_{\omega}$ is positive definite, and $\bar{\Omega}_{\omega}\succ 0$ also holds. Moreover, since $\Gamma_{m}$ has full row rank for every $m\in\mathcal{M}$ by its construction, it follows that $\Gamma_{m}\bar{\Omega}_{\omega}\Gamma_{m}^{\top}$ is positive definite. As $\bar{\omega}_{\ell}$ is Gaussian with covariance matrix $\bar{\Omega}_{\omega}$ and independent of $\hat{\chi}_{\ell}$, the state $\hat{\chi}_{\ell+1}$ conditioned on $\hat{\chi}_{\ell}=x$ and $\varphi(x)=m$ is Gaussian with mean $\Phi_{m}x$ and covariance matrix $\Gamma_{m}\bar{\Omega}_{\omega}\Gamma_{m}^{\top}\succ 0$. Therefore, the transition probability kernel $P$ admits a smooth density $P(x,dz)=p_{m}(x,z)\,dz$ for a fixed $x\in\mathbb{R}^{n_{x}}$ and $m=\varphi(x)$, where the mapping $z\mapsto p_{m}(x,z)$ is continuous and strictly positive on $\mathbb{R}^{n_{x}}$ [24, Sec. 4.4].
@@
The first equality is from (175). For the second equality, we note that $P(x,\cdot)=\nu_{\bar{\omega}}\circ\Psi_{x}^{-1}$ is absolutely continuous with respect to the Lebesgue measure and its Radon–Nikodym derivative is $p_{m}(x,\cdot)$. Hence, by the Radon–Nikodym theorem [31, Th. 6.10], it holds that
@@
for all $x\in\mathbb{R}^{n_{x}}$. By the Chapman–Kolmogorov equation with $P^{0}(x,\cdot)=\delta_{x}$ [24, Th. 3.4.2], for any $n\geq 1$, we obtain
@@
since $P(y,\mathcal{O})>0$ for all $y\in\mathbb{R}^{n_{x}}$ and $P^{n-1}(x,\cdot)$ is a probability measure on $\mathscr{B}(\mathbb{R}^{n_{x}})$[22, Ex. 1.4.1].
@@
where the first equality follows from the translation invariance of the Lebesgue measure [31, Th. 2.20]. Since $r\in L^{1}(\mathbb{R}^{n_{x}})$ and translations are continuous in $L^{1}$, it holds that
@@
by [31, Th. 9.5]. Hence, $\tilde{P}_{m}(x,\mathcal{A})$ is continuous in $x$ for each $m\in\mathcal{M}$. Since $\mathcal{M}$ is finite, the minimum of the finitely many continuous functions $x\mapsto\tilde{P}_{m}(x,A)$ is also continuous, which implies that $x\mapsto T(x,A)$ is continuous.
@@
Building upon Lemmas 7, 8, and 9, we are ready to prove that the process $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}_{\geq 0}}$ is $\tilde{f}_{m}$-ergodic. From Lemma 7, the process $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}{\geq 0}}$ is aperiodic and open set irreducible [24, Sec. 5.4, 6.1]. Furthermore, Lemma 8 shows that the chain is a T-chain. Hence, every compact set in $\mathbb{R}^{n_{x}}$ is petite [24, Th. 6.2.5], which yields that the set $\mathcal{D}$ is petite. Then, Lemmas 6 and 9 imply that the chain is non-evanescent [24, Th. 9.4.1], which is equivalent to Harris recurrence of $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}{\geq 0}}$ [24, Th. 9.2.2]. In addition, a Harris recurrent chain admits a unique invariant measure on $\mathscr{B}(\mathbb{R}^{n_{x}})$ [24, Th. 10.0.1]. Combining this fact with Lemma 9, we conclude that $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}{\geq 0}}$ is positive Harris recurrent with invariant measure $\pi_{\hat{\chi}}$ [24, Th. 11.0.1]. Moreover, it follows that $\tilde{f}_{m}(x)\geq 1$ for all $x\in\mathbb{R}^{n_{x}}$ by construction, which establishes the $\tilde{f}_{m}$-ergodicity of the chain $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}_{\geq 0}}$ [24, Th. 14.0.1]. ∎
@@
Since $P^{(1)}_{0}$ is positive definite by Lemma 4, from [23, Cor. 10.4.2], there exist constants $0<\underline{\lambda}\leq\overline{\lambda}<\infty$ such that

### v3

@@ I Introduction
This paper investigates a sparse control method that achieves desirable performance under intermittent actuation. Sparse signals, which remain zero over most of the time horizon, enable networked systems to achieve control objectives with limited resource usage and compact signal representations [1]. They can also reduce power consumption by saving actuation durations, which is particularly important in railways [2] and electric vehicles [3]. These advantages have motivated extensive research on the theoretical foundations of sparse control, including controllability [4], stability [5], and optimal control [6, 7]. Despite these advances, the fundamental challenge lies in reconciling the trade-off between closed-loop performance and input sparsity [8].
@@
Research on controller design under intermittent actuation has followed two main directions. One line of work studies optimal control problems with constraints on the number of control actions over the entire horizon [9, 10, 11]. Another approach incorporates sparsity-promoting regularization terms into the optimization problem rather than imposing explicit constraints [12, 13]. These studies provide near-optimal actuation strategies with theoretical performance bounds; however, their problem settings are restricted, such as considering noise-free systems [13, 11] or excluding control energy from the objective function [9, 10, 12]. For a more general setting, [14] has studied packetized dead-beat control with threshold-based event triggering. While their work quantifies the trade-off between communication cost and control performance, the actuation timings are determined by the predesigned threshold rule rather than being optimized. This limitation highlights the need for a principled method that jointly optimizes discrete actuation timings and continuous control laws, providing provable performance guarantees.
@@
This paper addresses these limitations by employing the rollout algorithm, a sequential optimization method in the context of dynamic programming [15, 16]. In a rollout scheme, control laws and scheduling decisions are optimized sequentially over a finite horizon, yielding suboptimal yet effective solutions. Within this framework, [17] has addressed quadratic cost minimization under an average communication rate constraint for piecewise-constant control inputs, where the constraint is imposed on the number of input value changes. Importantly, a theoretical analysis of the controller’s performance and stability has been provided. However, this setting differs from our objective, as we aim to promote sparse signals that remain identically zero over consecutive intervals rather than limiting input variations. Furthermore, [18] has studied the simultaneous minimization of a finite-horizon quadratic cost and the number of control actions. The authors proposed an algorithm that sequentially minimizes the one-step cost with respect to the control input and actuation timing, providing theoretical performance guarantees but no stability guarantees.
This paper develops a rollout-based framework for sparse intermittent actuation, directly addressing the trade-off between control performance and actuation rate. Using the rollout method, actuation timings are optimized in a receding horizon fashion, and the corresponding control inputs are determined online. Building on our preliminary work [18], which considered only a finite horizon with one-step optimization window and no stability analysis, we establish both performance and stability guarantees for a more general framework. While our analysis draws inspiration from [17], their results cannot be directly applied here, since the trade-off is formulated differently through the sparsity-promoting term. Finally, a numerical example demonstrates the effectiveness of the proposed method, showing advantages over periodic strategies and $\ell_{1}$-relaxed approximation of group sparsity.
@@
The remainder of this paper is organized as follows. In section II, we describe the problem formulation. Section III provides some preliminaries and presents the proposed algorithm, followed by a theoretical analysis of our framework in Section IV. Section V illustrates a numerical example, and Section VI concludes this paper.
@@ III Proposed Rollout-Based algorithm
This section presents algorithm design using a rollout approach. In Section III-A, we provide preliminaries on the rollout method. In section III-B, we derive the optimal periodic policy used as a basis for our algorithm, and Section III-C presents the proposed algorithm for Problem (14).
@@ III-C Algorithm Design
By the tower property of conditional expectation [21, Th. 5.1.6], we have
@@ IV Performance and Stability Guarantees of the Proposed Algorithm
Assumption 2(i) ensures that the steady-state Kalman gain $G$ has full row rank, since the covariance matrix $\tilde{\Sigma}$ is positive definite. Assumption 2(ii) guarantees that the minimum eigenvalue of $Q$ is strictly positive [22, Lemma 10.4.1]. When Assumption 2(iii) holds, $\Omega_{x_{0}}$ is positive definite since $(A,\Omega_{w}^{1/2})$ is controllable. Furthermore, this assumption implies $G_{k}=G$ for all $k$, i.e., the Kalman filter is stationary, which is the case of interest in this study.
@@
In (66), we write $\beta^{(m)}$ instead of $\beta^{(m)}_{k}$, since $\beta^{(m)}_{k}$ is independent of the time index $k$ under Assumption 2. By construction of Algorithm 1, we have $f_{m}(x)\geq 0$ for all $x\in\mathbb{R}^{n_{x}}$. Then, it follows that the Markov chain $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}_{\geq 0}}$ is ergodic, i.e., positive Harris recurrent [23, Ch. 14].
@@ VI Conclusion
We have proposed a sparse control framework that aims to balance the trade-off between control performance and actuation rate. We formulated the optimal control problem using actuation-rate regularization and developed a rollout-based algorithm to obtain a tractable solution. We established theoretical guarantees, showing that the proposed method achieves a performance guarantee with respect to periodic control and ensures mean-square stability of the closed-loop system. The effectiveness of the approach was demonstrated through a numerical example. Future work includes investigating performance guarantees for discounted-cost formulations and deriving strict performance improvement results over periodic control strategies.
@@
Assumption 1 and Lemma 3 ensure that $\tilde{P}_{p}$ is positive semidefinite at $\alpha=1$ [20, Sec. 8.4]. Assume that there exists a nonzero vector $\xi\in\mathbb{R}^{n_{x}}$ such that $\xi^{\top}\tilde{P}_{p}\xi=0$. Then, we consider the following optimal control problem for a noise-free system with the initial condition $x_{0}=\xi$:
@@
To establish the coerciveness of $V_{1}$, it is sufficient to show that the sublevel sets $\mathcal{L}_{r}(x)\coloneqq\{x\in\mathbb{R}^{n_{x}}\colon V_{1}(x)\leq r\}$ are precompact, i.e., their closures are compact, for each $r\in\mathbb{R}_{>0}$ [23, Sec. 9.4]. Since $P^{(1)}_{0}$ is positive definite by Lemmas 4 and 5, we have
@@
The set on the right-hand side is bounded for every $r\in\mathbb{R}_{>0}$, and hence $\mathcal{L}_{r}(x)$ is also bounded. Moreover, the continuity of $V_{1}$ implies that $\mathcal{L}_{r}(x)=V_{1}^{-1}([0,r])$ is closed [26, Th. 4.8]. By the Heine–Borel theorem, $\mathcal{L}_{r}(x)$ is compact [26, Th. 2.41], and thus its closure is also compact for each $r\in\mathbb{R}_{>0}$ [26, Th. 2.27], which establishes the claim. ∎
@@
Conditioning on $\hat{\chi}_{0}$ and applying both the tower property and the linearity of conditional expectation [21, Th. 5.1.2], the left-hand side of (125) becomes
@@
where $\pi_{\hat{\chi}}$ is a unique invariant measure of the chain $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}_{\geq 0}}$ [23, Th. 14.0.1]. Note that the existence of $\pi_{\hat{\chi}}$ is guaranteed by Theorem 2. The above integrand is strictly positive due to the open set irreducibility of $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}_{\geq 0}}$ (see Appendix -F for details). Applying the Stolz–Cesàro theorem [28] to (143) yields
@@
In the following, we show that $P$ is a transition probability kernel and that the process $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}_{\geq 0}}$ has the Markov property [29, Def. 8.4.3], i.e.,
@@
where $g_{m}(x)\coloneqq x^{\top}P^{(m)}_{0}x+\operatorname{tr}(P^{(m)}_{0}\Sigma)+\beta^{(m)}+\gamma^{(m)}$ and $h_{i,m}(x)\coloneqq g_{i}(x)-g_{m}(x)$. This implies that $\mathcal{C}_{m}$ is Borel measurable since each inverse image is open and Borel measurable [26, Th. 4.8]. Therefore, for any $m\in\mathcal{M}$, we have $\varphi^{-1}(\{m\})=\mathcal{C}_{m}\in\mathscr{B}(\mathbb{R}^{n_{x}})$, which establishes the measurability of $\varphi$ [30, Th. 1.12].
@@
Then, $\hat{\chi}_{\ell+1}=\Psi(\hat{\chi}_{\ell},\bar{\omega}_{\ell})$ for a given $\varphi$. For each $x$ and $\omega$, the mapping $(x,\omega)\mapsto\Phi_{m}x+\Gamma_{m}\omega$ is continuous, and hence measurable [30, Th. 1.7]. Since the indicator function is measurable, the mapping $(x,\omega)\mapsto\mathbb{I}_{\mathcal{C}_{m}}(x)(\Phi_{m}x+\Gamma_{m}\omega)$ is measurable. Therefore, $\Psi$ is also measurable for any $m\in\mathcal{M}$ [30, Th. 1.8].
@@
for later use. By construction, $\bar{\omega}_{\ell}$ is a measurable function of $(e_{\ell h},\,\bar{\mathcal{W}}_{\ell})$, and hence $\sigma(\bar{\omega}_{\ell})\subseteq\sigma(e_{\ell h})\vee\sigma(\bar{\mathcal{W}}_{\ell})$ holds from [30, Th. 1.12]. Moreover, the independence of the noise processes yields $\sigma(\bar{\mathcal{W}}_{\ell})\perp\sigma(\mathcal{I}_{\ell h})$ [31, Lemma 4.6]. Since $e_{\ell h}=x_{\ell h}-\mathbb{E}[x_{\ell h}\mid\mathcal{I}_{\ell h}]$ under Gaussian noises implies that $e_{\ell h}$ is independent of $\mathcal{I}_{\ell h}$ [31, Lemma 14.1], we have
@@
which implies $\sigma(\bar{\omega}_{\ell})\perp\sigma(\mathcal{I}_{\ell h})$ [31, Cor. 4.7]. Since $\mathscr{F}_{\ell}\subseteq\sigma(\mathcal{I}_{\ell h})$, we obtain
@@
by [21, Ex. 5.1.5], where $\nu_{\bar{\omega}}$ denotes the distribution of $\bar{\omega}_{\ell}$. For a given $x\in\mathbb{R}^{n_{x}}$, combining (155), (164), and (166) yields
@@
from (166) and (168) [21, Th. 5.1.5].
For a fixed $x\in\mathbb{R}^{n_{x}}$, we define the mapping $\Psi_{x}\colon\omega\mapsto\Psi(x,\omega)$. We note that $\Psi_{x}$ is measurable for each $x$ since $\Psi$ is measurable on $\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{x}h}$ [30, Th. 8.5]. Then, for every $\mathcal{A}\in\mathscr{B}(\mathbb{R}^{n_{x}})$, it holds that
@@
which means that the mapping $\mathcal{A}\mapsto P(x,\mathcal{A})$ is a pushforward measure induced by $\Psi_{x}$; hence, $\mathcal{A}\mapsto P(x,\mathcal{A})$ is a probability measure on $\mathscr{B}(\mathbb{R}^{n_{x}})$ [31, Lemma 1.24]. Furthermore, $x\mapsto P(x,\mathcal{A})$ is measurable for a fixed $\mathcal{A}$, which follows by combining (167) and Fubini’s theorem. Therefore, $P(x,\mathcal{A})$ satisfies the definition of the transition probability kernel [23, Sec. 3.4].
By the property of conditional expectation [29, Th. 5.4.6], for any sub $\sigma$-algebra $\mathscr{G}\subseteq\mathscr{B}(\mathbb{R}^{n_{x}})$, we have
@@
Since $\Omega_{v}\succ 0$ and Assumption 2 holds, the matrix $\Omega_{\omega}$ is positive definite, and $\bar{\Omega}_{\omega}\succ 0$ also holds. Moreover, since $\Gamma_{m}$ has full row rank for every $m\in\mathcal{M}$ by its construction, it follows that $\Gamma_{m}\bar{\Omega}_{\omega}\Gamma_{m}^{\top}$ is positive definite. As $\bar{\omega}_{\ell}$ is Gaussian with covariance matrix $\bar{\Omega}_{\omega}$ and independent of $\hat{\chi}_{\ell}$, the state $\hat{\chi}_{\ell+1}$ conditioned on $\hat{\chi}_{\ell}=x$ and $\varphi(x)=m$ is Gaussian with mean $\Phi_{m}x$ and covariance matrix $\Gamma_{m}\bar{\Omega}_{\omega}\Gamma_{m}^{\top}\succ 0$. Therefore, the transition probability kernel $P$ admits a smooth density $P(x,dz)=p_{m}(x,z)\,dz$ for a fixed $x\in\mathbb{R}^{n_{x}}$ and $m=\varphi(x)$, where the mapping $z\mapsto p_{m}(x,z)$ is continuous and strictly positive on $\mathbb{R}^{n_{x}}$ [23, Sec. 4.4].
@@
The first equality is from (175). For the second equality, we note that $P(x,\cdot)=\nu_{\bar{\omega}}\circ\Psi_{x}^{-1}$ is absolutely continuous with respect to the Lebesgue measure and its Radon–Nikodym derivative is $p_{m}(x,\cdot)$. Hence, by the Radon–Nikodym theorem [30, Th. 6.10], it holds that
@@
for all $x\in\mathbb{R}^{n_{x}}$. By the Chapman–Kolmogorov equation with $P^{0}(x,\cdot)=\delta_{x}$ [23, Th. 3.4.2], for any $n\geq 1$, we obtain
@@
since $P(y,\mathcal{O})>0$ for all $y\in\mathbb{R}^{n_{x}}$ and $P^{n-1}(x,\cdot)$ is a probability measure on $\mathscr{B}(\mathbb{R}^{n_{x}})$[21, Ex. 1.4.1].
@@
where the first equality follows from the translation invariance of the Lebesgue measure [30, Th. 2.20]. Since $r\in L^{1}(\mathbb{R}^{n_{x}})$ and translations are continuous in $L^{1}$, it holds that
@@
by [30, Th. 9.5]. Hence, $\tilde{P}_{m}(x,\mathcal{A})$ is continuous in $x$ for each $m\in\mathcal{M}$. Since $\mathcal{M}$ is finite, the minimum of the finitely many continuous functions $x\mapsto\tilde{P}_{m}(x,A)$ is also continuous, which implies that $x\mapsto T(x,A)$ is continuous.
@@
Building upon Lemmas 7, 8, and 9, we are ready to prove that the process $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}_{\geq 0}}$ is $\tilde{f}_{m}$-ergodic. From Lemma 7, the process $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}{\geq 0}}$ is aperiodic and open set irreducible [23, Sec. 5.4, 6.1]. Furthermore, Lemma 8 shows that the chain is a T-chain. Hence, every compact set in $\mathbb{R}^{n_{x}}$ is petite [23, Th. 6.2.5], which yields that the set $\mathcal{D}$ is petite. Then, Lemmas 6 and 9 imply that the chain is non-evanescent [23, Th. 9.4.1], which is equivalent to Harris recurrence of $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}{\geq 0}}$ [23, Th. 9.2.2]. In addition, a Harris recurrent chain admits a unique invariant measure on $\mathscr{B}(\mathbb{R}^{n_{x}})$ [23, Th. 10.0.1]. Combining this fact with Lemma 9, we conclude that $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}{\geq 0}}$ is positive Harris recurrent with invariant measure $\pi_{\hat{\chi}}$ [23, Th. 11.0.1]. Moreover, it follows that $\tilde{f}_{m}(x)\geq 1$ for all $x\in\mathbb{R}^{n_{x}}$ by construction, which establishes the $\tilde{f}_{m}$-ergodicity of the chain $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}_{\geq 0}}$ [23, Th. 14.0.1]. ∎
@@
Since $P^{(1)}_{0}$ is positive definite by Lemma 4, from [22, Cor. 10.4.2], there exist constants $0<\underline{\lambda}\leq\overline{\lambda}<\infty$ such that